A 20 gm ball is shot from a spring gun whose spring constant is 400 N/m.The spring is compressed by 5 cm. The maxmimum possible horizontal range of the ball for this compression is {g=10m/s}.

Grade:12th Pass

2 Answers

kathivarapu bhargav
10 years ago

1/2kx^2=1/2mv^2

solving we get v^2=50

max range=v^2/g=5m

SOURAV MISHRA
10 years ago

(1/2)kx2 = (1/2)mv2

the maximum range is given by v2/g = kx2/mg = 400*0.05*0.05/0.02*10 = 5m.